# Injectives obstruct Fourier-Mukai functors

### v1

@@ 7. Computing obstructions explicitly for Divisors
The observations from the previous section let us compute:
@@
$o_{i}\left(L\circ\iota\right)\in\mathrm{HH}^{i}\left(\mathrm{Inj},\mathrm{Inj}_{\eta}\right)_{i}$
@@
be an $\mathcal{A}_{i}$ functor from a $𝕜\Bbbk$-linear category $I$ then we have an isomorphism
@@
$o_{i}\left(f_{*}\circ\varphi\right)$ $\mapsto o_{i}\left(s_{*}\circ h_{*}\circ j\circ\varphi\circ\iota\right).$
@@
$\left(h\circ j\right)_{*}:\mathrm{HH}^{m}\left(\mathrm{Inj},f_{*}\mathrm{Inj}_{f_{*}\eta}\right)$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)$
$o_{m}\left(\varphi\right)$ $\mapsto o_{m}\left(h_{*}\circ j\circ\varphi\iota\right).$
For the compatibility with lifts observe that since $j$ and $h$ are homotopy equivalences on the image of the injectives and so admit quasi-inverses they are immediately compatible with intermediate lifts.
As by Proposition 6.7 the functor $s_{*}$ is the inclusion of the upper triangular matrices and by Proposition 6.8 an equivalence on the negative $\mathrm{Ext}$-groups. We get, since $\mathrm{Inj}$ is a $𝕜\Bbbk$-linear category
$s_{*}:\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)_{<0}$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\right)_{<0}$
@@
Let $f^{*}\dashv f_{*}:\mathcal{D}_{\infty}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\infty}\left(\mathrm{Qcoh}Y\right)$ be an adjuction of $\mathcal{A}_{\infty}$-functors and let $N$ be an $\mathcal{A}_{\infty}$-$I$-bimodule of the shape
$N:I\otimes I^{\mathrm{op}}$ $\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$
@@
for $I$ a $𝕜\Bbbk$-linear category. Then we have that the adjunction quasi-isomorphism
@@
induces a quasi-isomorphism of $\mathcal{A}_{\infty}$-$I$-bimodules.
$\begin{pmatrix}f_{*}N:I\otimes I^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f_{*}Ni,f_{*}Nj\right)\\ \alpha\otimes\beta\mapsto f_{*}N_{1}\left(\alpha\right)\otimes f_{*}N_{2}\left(\beta\right)\end{pmatrix}\cong\begin{pmatrix}I\otimes I^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}Ni,Nj\right)\\ \alpha\otimes\beta\mapsto f^{*}f_{*}N_{1}\left(\alpha\right)\otimes N_{2}\left(\beta\right)\end{pmatrix}.$
@@
$f^{*}\dashv f_{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right).$
@@
$\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
is a quasi-isomorphism of $\mathrm{dg}$-$I$-bimodules. By Definition of $\mathrm{dg}$-adjunction it is already a quasi-isomorphism of chain complexes of vectorspaces. In particular it suffices to prove that it is a morphism of $I$-bimodules.
@@
$f_{*}N_{1}\left(\alpha\right):f_{*}Ni^{\prime}\to f_{*}Ni$
, then we have the commutative diagram
@@
$\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(Y\right)\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
is compatible with both actions and induces a quasi-isomorphism of $\mathcal{A}_{\infty}$-$I$-bimodules. ∎
@@
The following is a rewritten proof to our setup of the well known fact that for a divisor $X\hookrightarrow Y$ the object $f_{*}\mathcal{F}\in\mathcal{D}_{\infty}\left(Y\right)$ usually is not formal over $Y$ for a sheaf $\mathcal{F}$, however, it is formal over $X$. In particular it is essential that we restrict in this proof to the bimodule structure over $\mathrm{Inj}$.
@@
$\cong\mathrm{Ext}^{*-2+m}_{X}\left(I\otimes C_{f}[1]\oplus I,J\otimes M\right)$ $f^{*}f_{*}\mathcal{O}_{X}\cong\otimes\left(\mathcal{O}_{X}\oplus C_{f}[1]\right)$
@@
on the $\mathrm{Inj}$-bimodule $\mathrm{H}^{<0}\mathrm{\mathrm{im}}f_{*}\circ L\circ\iota$. Now we compute in $\mathrm{HH}^{m}\left(\mathrm{Inj},\mathrm{H}^{2-m}\mathrm{\mathrm{im}}\psi_{\eta}\circ\iota\right)$ that the obstruction is independent of choices
@@
$\mathrm{Ext}_{\mathcal{X}_{\eta}}^{2-m}\left(L\left(I\right),L\left(J\right)\right)\cong\mathrm{Ext}_{\mathcal{Y}_{f_{*}\eta}}^{2-m}\left(f_{*}\circ L\left(I\right),f_{*}\circ L\left(J\right)\right)$
@@
The functor $\psi_{\eta}$ does not admit an $\mathcal{A}_{\infty}$/$\mathrm{dg}$-lift.
@@
By Lemma 7.11 we have that $o_{m}\left(\psi_{\eta}\circ\iota\right)$ does not vanish and is independent of choices. In particular $\psi_{\eta}\circ\iota$ cannot have an $\mathcal{A}_{\infty}$-lift. Since $\iota$ is dg, this means that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift. ∎
By [Kü22, Theorem 4.14] we can compute for a smooth hypersurface the kernels of $\mathrm{HH}^{*}\left(X,\mathcal{O}\left(p\right)\right)$ explicitly via twisted Hodge-diamonds, in particular all these kernels in $\mathrm{HH}^{\geq\dim X+3}\left(X,\mathcal{O}\left(p\right)\right)$ define non-$\mathrm{dg}$ functors.
@@
Finally, paired with [Kü22, Proposition 4.16] we get the following corollary

### v2

@@ 7. Computing obstructions explicitly for Divisors
The observations from the previous sections let us compute:
@@
$o_{i}\left(L\circ\iota\right)\in\mathrm{HH}^{i}\left(\mathrm{Inj},\mathrm{Inj}_{\eta}\right)_{i}.$
@@
be an $\mathcal{A}_{i}$ functor from a $𝕜\Bbbk$-linear category $\mathcal{I}$ then we have an isomorphism
@@
$o_{i}\left(f_{*}\circ\varphi\right)$ $\mapsto o_{i}\left(s_{*}\circ h_{*}\circ j\circ\varphi\right).$
@@
$\left(h\circ j\right)_{*}:\mathrm{HH}^{m}\left(\mathcal{I},f_{*}\mathrm{Inj}_{f_{*}\eta}\right)$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)$
$o_{m}\left(\varphi\right)$ $\mapsto o_{m}\left(h_{*}\circ j\circ\varphi\right).$
For the compatibility with lifts observe that since $j$ and $h$ are homotopy equivalences on the image of the injectives they admit quasi-inverses they are immediately compatible with intermediate lifts.
As by Proposition 6.7 the functor $s_{*}$ is the inclusion of the upper triangular matrices and by Proposition 6.8 an equivalence on the negative $\mathrm{Ext}$-groups. We get, since $\mathcal{I}$ is a $𝕜\Bbbk$-linear category that $s_{*}$ is an equivalence,
$s_{*}:\mathrm{HH}^{m}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)_{<0}$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\right)_{<0}$
@@
Let $f^{*}\dashv f_{*}:\mathcal{D}^{b}_{\infty}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\infty}^{b}\left(\mathrm{Qcoh}Y\right)$ be an adjuction of $\mathcal{A}_{\infty}$-functors and let $N$ be an $\mathcal{A}_{\infty}$$-\mathcal{I}$-bimodule of the shape
$N:\mathcal{I}\otimes\mathcal{I}^{\mathrm{op}}$ $\to\mathcal{D}^{b}_{\infty}\left(\mathrm{Qcoh}X\right)$
@@
for $\mathcal{I}$ a $𝕜\Bbbk$-linear category. Then we have that the adjunction quasi-isomorphism
@@
induces a quasi-isomorphism of $\mathcal{A}_{\infty}$$-\mathcal{I}$-bimodules.
$\begin{pmatrix}f_{*}N:\mathcal{I}\otimes\mathcal{I}^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f_{*}Ni,f_{*}Nj\right)\\ \alpha\otimes\beta\mapsto f_{*}N_{1}\left(\alpha\right)\otimes f_{*}N_{2}\left(\beta\right)\end{pmatrix}\cong\begin{pmatrix}\mathcal{I}\otimes\mathcal{I}^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}Ni,Nj\right)\\ \alpha\otimes\beta\mapsto f^{*}f_{*}N_{1}\left(\alpha\right)\otimes N_{2}\left(\beta\right)\end{pmatrix}.$
@@
$f^{*}\dashv f_{*}:\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right).$
@@
$\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
is a quasi-isomorphism of $\mathrm{dg}$$-\mathcal{I}$-bimodules. By definition of $\mathrm{dg}$-adjunction it is already a quasi-isomorphism of chain complexes of vectorspaces. In particular it suffices to prove that it is a morphism of $\mathcal{I}$-bimodules.
@@
$f_{*}N_{1}\left(\alpha\right):f_{*}Ni^{\prime}\to f_{*}Ni,$
then we have the commutative diagram
@@
$\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(Y\right)\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
is compatible with both actions and induces a quasi-isomorphism of $\mathcal{A}_{\infty}$$-\mathcal{I}$-bimodules. ∎
@@
The following is a rewritten proof to our setup of the well known fact that for a divisor $X\hookrightarrow Y$ and a sheaf $\mathcal{F}$ the object $f_{*}\mathcal{F}\in\mathcal{D}^{b}_{\mathrm{dg}}\left(Y\right)$ usually is not formal over $Y$, however, it is formal over $X$. In particular it is essential that we restrict in this proof to the bimodule structure over $\mathrm{Inj}$.
@@
$\cong\mathrm{Ext}^{*-2+m}_{X}\left(I\otimes C_{f}[1]\oplus I,J\otimes M\right)$ $\mathllap{f^{*}f_{*}\mathcal{O}_{X}\cong\otimes\left(\mathcal{O}_{X}\oplus C_{f}[1]\right)}$
@@
on the $\mathrm{Inj}$-bimodule $\mathrm{H}^{<0}\mathrm{\mathrm{im}}f_{*}\circ L\circ\iota$. Now we compute in
$\mathrm{HH}^{m}\left(\mathrm{Inj},\mathrm{H}^{2-m}\mathrm{\mathrm{im}}\psi_{\eta}\circ\iota\right)=\mathrm{HH}^{m}\left(\mathrm{Inj},\mathrm{H}^{2-m}\mathrm{\mathrm{im}}^{<0}\psi_{\eta}\circ\iota\right)$
that the obstruction is independent of choices
@@
$\mathrm{Ext}_{\mathcal{X}_{\eta}}^{2-m}\left(L\left(I\right),L\left(J\right)\right)\cong\mathrm{Ext}_{\mathcal{Y}_{f_{*}\eta}}^{2-m}\left(f_{*}\circ L\left(I\right),f_{*}\circ L\left(J\right)\right).$
@@
The functor
$\psi_{\eta}:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)$
does not admit an $\mathcal{A}_{\infty}$/$\mathrm{dg}$-lift.
@@
By Lemma 7.11 we have that $o_{m}\left(\psi_{\eta}\circ\iota\right)$ does not vanish and is independent of choices. In particular $\psi_{\eta}\circ\iota$ cannot have an $\mathcal{A}_{\infty}$-lift. Since $\iota$ is dg by Lemma 3.3, this means that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift. ∎
If the functor $\psi_{\eta}$ is continuous we can restrict our argument to the category of coherent sheaves. We conjecture that $\psi_{\eta}$ has this property as it is by definition defined on the subcategory of injectives and extended from there, see [RVdBN19]. However, as it is very ill-behaved we were not yet able to verify it.
Theorem 7.13.
Assume that $\psi_{\eta}$ is continuous. Then we have that the restriction
$\psi_{\eta}:\mathcal{D}^{b}\left(X\right)\to\mathcal{D}^{b}\left(Y\right)$
is not Fourier-Mukai.
Proof.
Assume that $\psi_{\eta}:\mathcal{D}^{b}\left(X\right)\to\mathcal{D}^{b}\left(Y\right)$ admits a dg-enhancement. Let $\widehat{\psi}_{\eta}:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)$ be the continuation of $\varphi_{\eta}$ to $\mathcal{D}^{b}\left(\mathrm{Qcoh}\_\right)$ via [Lur09, Proposition 5.3.5.10.], respectively [Toe07, Theorem 7.2], we have that this is the unique continuous extension of $\psi_{\eta}$. In particular we have, as $\psi_{\eta}$ was assumed to be continuous
$\widehat{\psi_{\eta}}=\psi_{\eta}::\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right).$
But as $\psi_{\eta}:\mathcal{D}^{b}\left(X\right)\to\mathcal{D}^{b}\left(Y\right)$ had a $\mathrm{dg}$-enhancement, $\widehat{\psi_{\eta}}$ would have a $\mathrm{dg}$-enhancement, which is contradiction to Theorem 7.12
This in particular means that the following corollaries could be restricted to the case of coherent sheaves if $\psi_{\eta}$ is continuous.
By [Kü24, Theorem 4.14] we can compute for a smooth hypersurface the kernels of $\mathrm{HH}^{*}\left(X,\mathcal{O}\left(p\right)\right)$ explicitly via twisted Hodge-diamonds, in particular all these kernels in $\mathrm{HH}^{\geq\dim X+3}\left(X,\mathcal{O}\left(p\right)\right)$ define non-$\mathrm{dg}$ functors.
@@
Finally, paired with [Kü24, Proposition 4.16] we get the following corollary